OpenStudy (candyme):

What method would be best to solve this system of equations? (Addition, Subtraction, Subsitution) How would you solve it? Please include steps, because I want to learn how to do this.

OpenStudy (candyme):

\[2x+4y=8\]\[y-3=x\]

OpenStudy (anonymous):

Substitute x=y-3 into the top equation, i.e. Change 2x in the top eq to 2(y-3)

OpenStudy (anonymous):

y-3=x or x=y-3 -----(1) put the value of x in eq (2) i.e, 2x+4y=8 2(y-3) +4y=8 2y-6+4y=8 2y+4y=8+6 6y=14 y=14/6=7/3 put the value of y in eq (1) x=y-3 x=7/3-3 x=7-9/3 x=-2/3

OpenStudy (candyme):

Thanks to you both, this really helped me see what I should do!